OpenStudy (thegrimreaper666):

What is the length of HD(with a line on top) when BH=20 inches, BC=12 inches, and DC= 14 inches? Round the answer to the nearest tenth of an inch.

OpenStudy (mathstudent55):

Do you know the Pythagoras theorem? \(a^2 + b^2 = c^2\)

OpenStudy (thegrimreaper666):

I do.

OpenStudy (mathstudent55):

Use the Pythagoras theorem on triangle BCH to find CH. Then use the Pythagoras theorem again on triangle CDH to find HD.

OpenStudy (thegrimreaper666):

So, 400-144=256

OpenStudy (thegrimreaper666):

And the square root of 256 is 16

OpenStudy (mathstudent55):